# KITchen: A Real-World Benchmark and Dataset for 6D Object Pose Estimation in Kitchen Environments

### v2

@@ Abstract
Despite the recent progress on 6D object pose estimation methods for robotic grasping, a substantial performance gap persists between the capabilities of these methods on existing datasets and their efficacy in real-world grasping and mobile manipulation tasks, particularly when robots rely solely on their monocular egocentric field of view (FOV). Existing real-world datasets primarily focus on table-top grasping scenarios, where a robot arm is placed in a fixed position and the objects are centralized within the FOV of fixed external camera(s). Assessing performance on such datasets may not accurately reflect the challenges encountered in everyday grasping and mobile manipulation tasks within kitchen environments such as retrieving objects from higher shelves, sinks, dishwashers, ovens, refrigerators, or microwaves. To address this gap, we present KITchen, a novel benchmark designed specifically for estimating the 6D poses of objects located in diverse positions within kitchen settings. For this purpose, we recorded a comprehensive dataset comprising around 205k real-world RGBD images for 111 kitchen objects captured in two distinct kitchens, utilizing a humanoid robot with its egocentric perspectives. Subsequently, we developed a semi-automated annotation pipeline, to streamline the labeling process of such datasets, resulting in the generation of 2D object labels, 2D object segmentation masks, and 6D object poses with minimal human effort. The benchmark, the dataset, and the annotation pipeline will be publicly available at https://kitchen-dataset.github.io/KITchen.
@@ II-A Objects Datasets
The current landscape of 6D pose estimation methods is diverse, ranging from traditional techniques such as template matching [27, 20, 42, 36] and correspondences with locally invariant features [13, 12, 38] to the current advanced deep learning SOTA render & compare approaches [32, 49]. These approaches provide the 6D poses of novel objects by rendering many views of the object during inference using its 3D CAD model and then passing these rendered views with the received cropped image of the object obtained by any 2D object detectors [43, 34, 53, 33, 47] to a coarse model which classifies which rendered image best matches the input image. Finally, they pass the initial pose to a refiner network to estimate an updated 6D pose of the object. In this work, we leverage MegaPose [32], Segment Anything [30], and YOLOv5 [26] to annotate our dataset.
@@ III-C Annotation Pipeline
Annotating objects with their ground truth 6D poses is a labor-intensive and time-consuming task. Although some of the recently published datasets attempted to semi-automate the annotation process. For instance, GraspNet-1Billion’s approach [18] relies on manually annotating the first frame of each scene, then leveraging recorded camera poses to calculate the objects’ poses in the following frames. However, this method was not optimal for our dataset, as KITchen has many more diverse scenes per kitchen compared to the simple setup used in GraspNet-1Billion, resulting in significantly more effort required for manual annotation. Another attempt to semi-automate the annotation process was presented by HANDAL [19], but their approach assumes a single object in the scene, making it unsuitable for our dataset, which contains multiple objects $10-50$ objects in each scene. To overcome the limitations of existing approaches and streamline the annotation process, we propose a semi-automated annotation pipeline. This pipeline generates three types of annotations: 2D object bounding boxes, 2D segmentation masks, and 6D poses, see 6.
Figure 6: Our proposed annotation pipeline. The pipeline starts 3D meshes of dataset objects as input, which are processed by BlenderProc2 to generate synthetic data with 2D bounding boxes. This annotated 2D data is used to train a YOLOv5 2D object detector. Subsequently, real-world recorded data is fed into the trained model, and the output is manually inspected for correct and incorrect labeling. The correctly labeled images are used for model refinement, which is then validated on the incorrectly labeled images. This iterative process continues until all images are correctly labeled. The correctly labeled images are then passed to Segment Anything (SAM) to generate masks. Finally, the images, along with the 2D labels and 3D meshes, are fed into MegaPose to generate 6D poses for detected objects. Manual inspection of poses is performed using contour and mesh overlay images, and corrected annotations are used to iteratively fine-tune MegaPose until the entire dataset is accurately annotated.
@@
The pipeline starts by receiving the collected 3D CAD object models for the dataset, then it generates around 100K annotated photo-realistic synthetic RGBD images with 2D bounding boxes using BlenderProc2 [15]. These synthetic images are used to finetune a pretrained YOLOv5 model [26] for 2D object detection. Subsequently, the trained model is applied to our real-world data, and manually classified images are inspected to distinguish correctly labeled ones. The model is then fine-tuned iteratively until all real-world data is accurately labeled with 2D object labels.
@@
To generate the 6D poses for the objects in the images, we pass the 2D bounding boxes which are generated using the fine-tuned YOLOv5 object detection model alongside the 3D CAD models of the detected objects with the input image into MegaPose [32]. The output 6D poses are used to overlay contours and meshes on the images for manual inspection. The MegaPose model is fine-tuned with corrected labeled data iteratively until all data are accurately annotated. The entire annotation pipeline is illustrated in Fig. 6 and several illustrative examples of the output of each step are demonstrated in Fig. 7.
@@ IV-A Problem Statement
The benchmark is designed to address the object 6D pose estimation problem, where the model receives an image $I$ from the dataset $D$, where $D$ is a set of RGBD images. The image $I$ contains a set of objects $\{{o\}^{n}_{i=0}}$. The model has access to the $M$, where $M$ is a set of 3D meshes of all objects $O$ in the dataset $D$. The objective is to estimate the pose $P$ of all objects $\{{o\}^{n}_{i=0}}$ in each image $I$, where $P=[R,T;0,1]$, where $R$ is a $3\times 3$ rotational matrix that describes the rotation of each of $\{{o\}^{n}_{i=0}}$ to the robot camera’s frame and $T$ is the translation vector to the origin of robot camera’s coordinate system.

### v3

@@ Abstract
Despite the recent progress on 6D object pose estimation methods for robotic grasping, a substantial performance gap persists between the capabilities of these methods on existing datasets and their efficacy in real-world grasping and mobile manipulation tasks, particularly when robots rely solely on their monocular egocentric field of view (FOV). Existing real-world datasets primarily focus on table-top grasping scenarios, where a robot arm is placed in a fixed position and the objects are centralized within the FOV of fixed external camera(s). Assessing performance on such datasets may not accurately reflect the challenges encountered in everyday grasping and mobile manipulation tasks within kitchen environments such as retrieving objects from higher shelves, sinks, dishwashers, ovens, refrigerators, or microwaves. To address this gap, we present KITchen, a novel benchmark designed specifically for estimating the 6D poses of objects located in diverse positions within kitchen settings. For this purpose, we recorded a comprehensive dataset comprising around 205k real-world RGBD images for 111 kitchen objects captured in two distinct kitchens, utilizing a humanoid robot with its egocentric perspectives. Subsequently, we developed a semi-automated annotation pipeline, to streamline the labeling process of such datasets, resulting in the generation of 2D object labels, 2D object segmentation masks, and 6D object poses with minimal human effort. The benchmark, the dataset, and the annotation pipeline are publicly available at https://kitchen-dataset.github.io/KITchen.
@@ II-A Objects Datasets
The current landscape of 6D pose estimation methods is diverse, ranging from traditional techniques such as template matching [34, 35, 36, 37] and correspondences with locally invariant features [38, 39, 40] to the current advanced deep learning SOTA render & compare approaches [13, 41]. These approaches provide the 6D poses of novel objects by rendering many views of the object during inference using its 3D CAD model and then passing these rendered views with the received cropped image of the object obtained by any 2D object detectors [42, 43, 44, 45, 46] to a coarse model which classifies which rendered image best matches the input image. Finally, they pass the initial pose to a refiner network to estimate an updated 6D pose of the object. In this work, we leverage MegaPose [13], Segment Anything [47], and YOLOv8 [48] to annotate our dataset.
@@ III-C Annotation Pipeline
Annotating objects with their ground truth 6D poses is a labor-intensive and time-consuming task. Although some of the recently published datasets attempted to semi-automate the annotation process. For instance, GraspNet-1Billion’s approach [31] relies on manually annotating the first frame of each scene, then leveraging recorded camera poses to calculate the objects’ poses in the following frames. However, this method was not optimal for our dataset, as KITchen has many more diverse scenes per kitchen compared to the simple setup used in GraspNet-1Billion, resulting in significantly more effort required for manual annotation. Another attempt to semi-automate the annotation process was presented by HANDAL [51], but their approach assumes a single object in the scene, making it unsuitable for our dataset, which contains $10-50$ objects in each scene. To overcome the limitations of existing approaches and streamline the annotation process, we propose a semi-automated annotation pipeline. This pipeline generates three types of annotations: 2D object bounding boxes, 2D segmentation masks, and 6D poses, see Fig. 6.
Figure 6: Our proposed annotation pipeline. The pipeline starts with 3D meshes of dataset objects as input, which are processed by BlenderProc2 to generate synthetic data with 2D bounding boxes. This annotated 2D data is used to train a YOLOv8 2D object detector. Subsequently, real-world recorded data is fed into the trained model, and the output is manually inspected for correct and incorrect labeling. The correctly labeled images are used for model refinement, which is then validated on the incorrectly labeled images. This iterative process continues until all images are correctly labeled. The correctly labeled images are then passed to Segment Anything (SAM) to generate masks. Finally, the images, along with the 2D labels and 3D meshes, are fed into MegaPose to generate 6D poses for detected objects. Manual inspection of poses is performed using contour and mesh overlay images, and corrected annotations are used to iteratively fine-tune MegaPose until the entire dataset is accurately annotated.
@@
The pipeline starts by receiving the collected 3D CAD object models for the dataset, then it generates around 100K annotated photo-realistic synthetic RGBD images with 2D bounding boxes using BlenderProc2 [52]. These synthetic images are used to finetune a pretrained YOLOv8 model [48] for 2D object detection. Subsequently, the trained model is applied to our real-world data, and manually classified images are inspected to distinguish correctly labeled ones. The model is then fine-tuned iteratively until all real-world data is accurately labeled with 2D object labels.
@@
To generate the 6D poses for the objects in the images, we pass the 2D bounding boxes which are generated using the fine-tuned YOLOv8 object detection model alongside the 3D CAD models of the detected objects with the input image into MegaPose [13]. The output 6D poses are used to overlay contours and meshes on the images for manual inspection. The MegaPose model is fine-tuned with corrected labeled data iteratively until all data are accurately annotated. The entire annotation pipeline is illustrated in Fig. 6 and several illustrative examples of the output of each step are demonstrated in Fig. 7.
@@ IV-A Problem Statement
The benchmark is designed to address the object 6D pose estimation problem, where the model receives an image $I$ from the dataset $D$, where $D$ is a set of RGBD images. The image $I$ contains a set of objects $\{{o\}^{n}_{i=0}}$. The model has access to the $M$, where $M$ is a set of 3D meshes of all objects $O$ in the dataset $D$. The objective is to estimate the pose $P$ of all objects $\{{o\}^{n}_{i=0}}$ in each image $I$, where $P=[R,T;0,1]$, where $R$ is a $3\times 3$ rotational matrix that describes the rotation of each of detected objects $\{{o\}^{n}_{i=0}}$ to the robot camera’s frame and $T$ is the translation vector to the origin of robot camera’s coordinate system.